OpenStudy (anonymous):

What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

An arithmetic sequence has a general form like this, \[a_n=a_1+d\cdot(n-1)\]with d, the distance between two consecutive terms. You can deduce from the problem that, d=-6, and that a_11=12 when n is equal to 11. Can you follow from this point?

OpenStudy (anonymous):

The answer choices are: a: -174 b: -176 c: -180 d: -186 e: -240 The answer I got wasnt on the list. so i probably am really confused with this

OpenStudy (john_es):

The first thing you need is a_1. In order to obtain it let's introduce the data of the problem \[a_{11}=12=a_1+(-6)(11-1)\Rightarrow a_1=72\] Do you understand this step?

OpenStudy (anonymous):

Yes.

OpenStudy (john_es):

Now, you can calculate a_43, the only thing you need to do is put n=43 in the general form, \[a_{43}=72+(-6)(43-1)=\] Do it.

OpenStudy (anonymous):

-180

OpenStudy (john_es):

Exactly.

OpenStudy (anonymous):

Ohhh. I got it. Thanks so much!!

OpenStudy (john_es):

You're welcome.